# SKVQ: Sliding-window Key and Value Cache Quantization for Large Language Models

### v2

@@ Abstract
Large language models (LLMs) can now handle longer sequences of tokens, enabling complex tasks like book understanding and generating lengthy novels. However, the key-value (KV) cache required for LLMs consumes substantial memory as context length increasing, becoming the bottleneck for deployment. In this paper, we present a strategy called SKVQ, which stands for sliding-window KV cache quantization, to address the issue of extremely low bitwidth KV cache quantization. To achieve this, SKVQ rearranges the channels of the KV cache in order to improve the similarity of channels in quantization groups, and applies clipped dynamic quantization at the group level. Additionally, SKVQ ensures that the most recent window tokens in the KV cache are preserved with high precision. This helps maintain the accuracy of a small but important portion of the KV cache. SKVQ achieves high compression ratios while maintaining accuracy. Our evaluation on LLMs demonstrates that SKVQ surpasses previous quantization approaches, allowing for quantization of the KV cache to 2-bit keys and 1.5-bit values with minimal loss of accuracy. With SKVQ, it is possible to process context lengths of up to 1M on an 80GB memory GPU for a 7b model and up to 7 times faster decoding. Code will be released at https://github.com/cat538/SKVQ.
@@ 1 Introduction
Recently, large Language Models (LLMs) have achieved great success in the area of artificial intelligence. With the advancement of LLMs, the need to support the longer context has grown. For instance, OpenAI GPT-4 Turbo can handle 128k tokens (Achiam et al., 2023), and Google Gemini 1.5 can process up to 1 million tokens (Team et al., 2023). This expanded token support enables LLMs to tackle more complex tasks like book reading, large image understanding, and video processing, making them more versatile. LLM inference operates in an auto-regressive manner, generating sentences token by token. To reduce the computation overhead, inference system always store the key and value activations in memory and reuse them during subsequent token generation steps. The saved data is known as key and value cache (KV Cache). With the increasing popularity of utilizing LLM for long sequence tasks, the KV cache consumes a significant amount of memory. On the other hand, the large amount of KV cache can also bring a large amount of memory access in the attention mechanism when generating the output tokens. The system will be stuck on the memory access, known as the memory-bound problem in LLM inference (Yuan et al., 2024).
@@
To evaluate the effectiveness of our method, we experiments on models of LLaMA(Touvron et al., 2023) and Mistral(Jiang et al., 2023) family. The experiments show that our methods can quantize the key cache into 2 bits and value cache into 1.5 bits with almost no accuracy drop. Compared with the previous quantization method, our approach can achieve optimal performance under different average bit widths as shown in Figure 1. Our performance analysis shows SKVQ enables 1M context length in a single A100-80GB for a 7b model. As for the inference latency, in the case of batch size 128 and sequence length 200k, the theoretical 7x speedup in decoding phase can be achieved  [^1] The performance analysis can be found in Appendix C.
@@ 3.1 Clipped Dynamic Quantization with Channel Reorder
To address this problem, some methods have proposed using additional quantization parameters or keeping certain channels in float format to handle outliers (Dettmers et al., 2022). However, we have noticed that the concept of outliers is relative. The channels with the highest values are outliers compared to the medium-sized channels, and the medium-sized channels are outliers compared to the small-sized channels. Other methods propose smoothing the difference between channels by multiplying an extra factor before quantization (Shao et al., 2023; Yue et al., 2024). However, these methods do not take into account the differences in token dimensions. The magnitude of values can vary between different tokens. We have observed that the variation in magnitude of non-outlier channels is relatively high. Specifically, some channels experience magnitude changes of several times or even dozens of times. Smoothing is not effective in addressing this phenomenon, especially in extremely low bitwidth quantization.
@@
where $P_{k}\in\mathbb{R}^{C_{in}\times C_{in}}$ and $P_{v}\in\mathbb{R}^{C_{in}\times C_{in}}$ are channel reorder matrix of Key and Value respectively  [^2] We fuse the channel reorder index into the the projection weight matrix of attention module. We describe the fusion in the AppendixA.. In our algorithm, the index is calculated based on the statistical characteristics of each channel. Specifically, we extract the distribution feature of each channel and then use the KMeans algorithm to cluster channels with similar characteristics into the same group. We also compared channel reordering with mathematical equivalent smoothing, and the results in AppendixD demonstrated the effectiveness of the former.
@@ 3.2 Sliding Window Quantization Strategy
Locality. Many previous works have shown that attention module has very strong locality(Kovaleva et al., 2019; Beltagy et al., 2020; Ge et al., 2023). That means at each time step, the attention module pays more attention to the recently generated tokens. We indicate that compared to the large amount of but less important content in the previous KV cache , protecting the accuracy of these small portion of but more important caches in KV cache quantization is critical.
@@
Motivated by this, we proposed a sliding window quantization strategy, which retain the latest KV cache of a window $w$ tokens to high precision. The workflow is shown in Figure 3. 1) In prefill phase, for each transformer block, after the KV cache is generated, we first compute attention with full precision KV cache , then quantize the KV cache with the last window_size token cache pairs reserved as full-precision. 2) In decode phase, we only process the token which slides out of the window at each time step. This approach ensures the KV cache of each transformer block generated in the prefill phase are lossless. It also enhance the generated content quality by utlizing the locality of attention module in the decode phase.
@@
Important KV Cache Filter. Except the recent generated tokens, there are some tokens that are sensitive to quantization. We also explored other method to identify important tokens that their KV cache should be kept in high precision. Inspired by (Xiao et al., 2023), the first few tokens of prompt are also very important for the whole generation process, so we add attention sink to the filter rules, i.e., the first few tokens are reserved to high precision. We observed it is effective to keep a small number of sink tokens high precision. Since the positions of sink tokens are fixed, it is easy to implement and we enable it in our experiments. Some cache eviction method monitor for each token its cumulative sum of attention score, then treat these scores as token frequency and only keep the most frequent tokens in the KV cache (Liu et al., 2023; Zhang et al., 2023), which are called heavy hitters. A straightforward idea is to keep heavy hitters to high precision. However, we did not enable it in our experiments for two reasons: 1) The improvement on prediction accuracy by keeping heavy hitters high precision is not significant. 2) If FlashAttention(Dao, 2023) is used, we can’t directly obtain the attention score. It is unfriendly to implement in existing inference framework. We believe that there are better methods available to identify important KV caches. Therefore, we have kept this as an interface (filter rules in Figure 3) in our implementation, allowing for the addition of new filters in future research.
@@
By retaining small portion of tokens to high precision, we obtain a substantial performance gain in the long context task, while at the same time incurring almost no additional overhead. We will show the impact of window size on quantization performance in Section 4.3.
@@ 4.2 Main Results and Analysis
Model Method LCC RepoBench-P PR-en TREC 2wikimqa GovReport MQA-zh Average
@@
LongBench Results. The performance of SKVQ in the LongBench datasets is summarised in Table 1. We compare our method with Smoothquant(Xiao et al., 2022), RPTQ(Yuan et al., 2023) KIVI(Liu et al., 2024) and per-token RTN(Round To Nearest). Smoothquant and RPTQ are LLM weight-activation quantization schemes. We use them to quantize KV cache without involving model weights and other activation. KIVI(Liu et al., 2024) is a recent 2-bit asymmetric quantization scheme specially designed for KV cache . We set the group size of all the methods to 128. SKVQ utilizes reordering which leads to unequal sizes for each group. In order to ensure the fairness of the comparison, we control the number of groups in SKVQ to ensure the average group size is 128. The window size in SKVQ is set to 128 and the residual length in KIVI is set to 128. $\alpha$ in Smoothquant is set to 1.0 to make the smooth transformation completely inclined to KV cache . Table 1 suggests that SKVQ is an effective method for KV cache compression that outperforms previous quantization approaches across various hard long context generation tasks. We also evaluate Vicuna-v1.5-7b-16k and LongChat-v1.5-7b-32k, the results is in AppendixE.
@@
These results demonstrate that it is practical to quantize the key-value cache into extremely low-bitwidth for these tasks. More result on needle in haystack test can be found in Appendix B.
@@ 4.3 Ablation Study
Breakdown of different components of SKVQ. We study the accuracy gain or loss of different quantization techniques used in SKVQ. We first use RTN and adopt per-token quantization with group size 32. We then apply other quantization techniques used in SKVQ, i.e. sliding window, clipping, channel reorder, attention sink and FP8. We present the LongBench average score in Table 3. Attention sink size is set to 5, i.e. the first 5 token cache pairs are retained to full-precision. We see that sliding-window and channel reorder can significantly enhance the accuracy. FP8(E4M3) stands for using FP8 datatype to store the per-group quantization parameters, i.e. scale and zero-point. In our study, Using FP8 will results in a relative minor accuracy decreasing compared with FP16. However, at extremely low-bitwidth and fine-grained group size, using FP8 to store the quantization parameter significantly reduce storage overhead. For example, KV 2bits with group size 32, if we use FP16 to store quantization parameters, the average bits of element in KV cache is $2+16*2/32=3$, but if we use FP8, then average bits is $2+8*2/32=2.5$, which is 50% smaller.
@@ 5 Conclusion
Appendix A Detailed Implementations
We describe our algorithm as shown in algorithm1. The subroutine $\mathrm{get\_permutation\_matrix}$ and $\mathrm{get\_group\_clipping}$ is described in sec3.1. It’s worth noting that the prologue only needs to be executed once before deploying, and we do not pay for it during the inference phase.
SKVQ Parameter: window size $W$, group size $G$, filter rules $F\leftarrow\{f_{1},f_{2},\cdots\}$, processed KV cache length $processed\leftarrow 0$
Attention Module Parameter: $W_{q},W_{k},W_{o}\in\mathbb{R}^{d\times d}$
Prologue :
       $P_{k},P_{v},group\_indices\leftarrow\mathrm{get\_permutation\_matrix}(calibration\_set)$
       $clipping\leftarrow\mathrm{get\_group\_clipping}(calibration\_set,P,group\_indices)$
       $W_{k}\leftarrow P_{k}\cdot W_{k}$
       $W_{v}\leftarrow P_{v}\cdot W_{v}$
end Prologue
Input: $\mathbf{X}\in{\mathbb{R}^{l\times d}}$, $\mathbf{K_{cache}},\mathbf{V_{cache}}\in\mathbb{R}^{h\times d},~{}\mathrm{where}~{}h~{}$ is context length(prefill phase $h=0$), $l$ is current input length(prefill phase $l=\mathrm{len}(prompt)$, decode phase $l=1$)
Algorithm algo(Attention module with SKVQ algorithm):
       $Q=\mathbf{X}\cdot W_{q}$, $K=\mathbf{X}\cdot W_{k}$, $V=\mathbf{X}\cdot W_{v}$
       $\mathbf{K_{cache}}\leftarrow\mathrm{dequant}(\mathbf{K_{cache}})$
       $\mathbf{V_{cache}}\leftarrow\mathrm{dequant}(\mathbf{V_{cache}})$
       $\mathbf{K_{cache}}\leftarrow\mathrm{concat}(\mathbf{K_{cache}},K)$
       $\mathbf{V_{cache}}\leftarrow\mathrm{concat}(\mathbf{V_{cache}},V)$
       $S\leftarrow Q\cdot\mathrm{reorder}(\mathbf{K_{cache}})^{T}$
       $O\leftarrow S\cdot\mathrm{reorder}(\mathbf{V_{cache}})\cdot W_{o}$
       $ctx\_len\leftarrow\mathrm{len}(\mathbf{V_{cache}})$
       $indices\leftarrow[processed:ctx\_len-W$]
       if $indices\neq\emptyset$ then
             $kmask\leftarrow[processed:ctx\_len-W;False]$
             $vmask\leftarrow[processed:ctx\_len-W;False]$
             for $filter~{}~{}\mathrm{in}~{}~{}F$ do
                   $kmask\leftarrow filter(\mathbf{K_{cache}}[indices])\wedge kmask$
                   $vmask\leftarrow filter(\mathbf{V_{cache}}[indices])\wedge vmask$
             end for
            $\mathbf{K_{cache}}[indices]\leftarrow\mathrm{clipping\_quant}(\mathbf{K_{cache}}[indices],kmask)$
             $\mathbf{V_{cache}}[indices]\leftarrow\mathrm{clipping\_quant}(\mathbf{V_{cache}}[indices],vmask)$
             $processed\mathrel{+}=\mathrm{len}(indices)$
       end if
      return $\mathbf{O}$
end Algorithm
function $\mathrm{clipping\_quant}(X,mask)$:
       $groups\leftarrow\mathrm{split\_groups}(X,group\_indices)$
       $quant\_cache\leftarrow\emptyset$ for $group~{}~{}\mathrm{in}~{}~{}groups$ do
             $group\_min,group\_max\leftarrow\mathrm{minmax}(gruop)$
             $group\_min\leftarrow clipping[group]\times group\_min$
             $group\_max\leftarrow clipping[group]\times group\_max$
             $quant\_group\leftarrow\mathrm{group\_quant(group\_min,group\_max,group)}$
             $quant\_group[mask]\leftarrow group[indices]$
             $quant\_cache\leftarrow\mathrm{concat}(quant\_cache,quant\_group)$
       end for
      return quant_cache
end function
Algorithm 1 SKVQ Algorithm
Appendix B More Experimental Results of Needle in Haystack
Figure 7: Comparison of SKVQ with KIVI on 32k context length needle in haystack test. The baseline score is 268.5. We vary the group size from 64 to 128, and vary the quantization bits from (key 2bits, value 2bits) to (key 2bits, value 1.5bits).
The results in Figure7 show SKVQ is clearly better than KIVI, especially in K2V1.5-g128, SKVQ achieve the same level with FP16, while KIVI suffers significant accuracy loss. These results shows the robustness of our approach.
Appendix C Memory and Latency Analysis
In order to further illustrate the benefits of quantizing KV cache to extremely low-bitwidth, we use LLM-Viewer(Yuan et al., 2024) to analyze the benefits in terms of memory consumption and inference latency. The result is shown in table5. When batch size and sequence length are relatively large, KV cache dominates almost all the memory consumption, and load KV cache becomes the performance bottleneck of the entire inference system. By quantizing KV cache with SKVQ, we can significantly reduce both latency and memory consumption. The analysis result shows SKVQ enables 1M context length in a single A100-80GB. As for the inference latency, we show the results of decoding phase, which domain the inference time in long context tasks. In the case of batch size 128 and sequence length 200k, the theoretical 7x speedup can be achieved.
Batch Size Seq Length Latency(ms) / Memory(GB) FP16 KV4 KV2
1 32k Inference Time 10.6 7.5 7
Memory Access 21.6 15.3 14.3
Memory Consumption 29.7 17.2 15.1
128k Inference Time 23.1 10.8 8.7
Memory Access 47.2 22 17.8
Memory Consumption 80.1 29.7 21.4
200k Inference Time 32.5 13.3 10
Memory Access 66.3 27 20.5
Memory Consumption 118 39.2 26.1
64 32k Inference Time 274.1 76.6 43.7
Memory Access 559 156 89.1
Memory Consumption 1100 282 147
128k Inference Time 1100 286.4 154.8
Memory Access 2200 584 316
Memory Consumption 4300 1100 551
200k Inference Time 1700 443 238.1
Memory Access 3400 905 485
Memory Consumption 6700 1700 853
128 32k Inference Time 541.8 146.8 81
Memory Access 1100 299 165
Memory Consumption 2200 550 282
128k Inference Time 2100 566.4 303.1
Memory Access 4400 1200 618
Memory Consumption 8600 2200 1100
200k Inference Time 3300 881.1 469.7
Memory Access 6800 1800 958
Memory Consumption 13400 3400 1700
Table 5: LLaMA-7B memory and latency analysis with roof line model.The hardware platform is A100 80G, we assume flash-attention is used.
Appendix D Comparison Between Smooth and Reorder
To improve the accuracy of per-token quantization, we utilize the reorder to cluster similar channels together. There are also other methods to improve the accuracy, one of them is smoothing, which is adopted in (Xiao et al., 2022; Shao et al., 2023; Yue et al., 2024). This approach smooth the difference between channels by multiplying an extra factor before quantization. We explore and compare smoothing with reordering, the experimental results are shown in table6. SKVQ-smooth represents our sliding window strategy together with smoothing and SKVQ-reorder represents the approach we described in 3.1. The results demonstrate that reordering can effectively improve the per-token quantization performance while the smoothing cannot. This is mainly because smoothing does not take into account the differences in token dimensions.
Model Method LCC RepoBench-P PR-en TREC 2wikimqa GovReport MQA-zh Average
LLaMA-2-7B-chat FP16 52.33 44.05 10.25 63 32.09 27.29 11.39 38.50
SKVQ-reorder 50.69 45.4 5.5 63 28.5 27.07 10.7 37.50
SKVQ-smooth 48.93 40.12 4.75 62.5 26.75 23.19 7.93 34.77
LLaMA-2-13B-chat FP16 50.54 52.1 15.25 68.5 13.21 27.52 7.23 38.83
SKVQ-reorder 49.53 49.76 12.25 67.5 14.03 26.68 6.63 37.53
SKVQ-smooth 47.78 47.28 7.5 67 11.61 24.07 5.55 35.34
Mistral-7B FP16 68.06 60.46 17.71 68 10.87 20.09 17.1 45.51
SKVQ-reorder 67.81 60.54 13.21 67 10.91 17.72 15.9 43.47
SKVQ-smooth 64.18 57.95 9.49 63.5 10.11 13.99 12.77 41.52
Mistral-7B-Instruct FP16 55.07 48.96 60 70 22.63 31.18 42.74 48.66
SKVQ-reorder 54.86 49.05 56.42 70 20.94 30.82 42.4 46.23
SKVQ-smooth 49.83 45.74 40.42 66 17.11 28.32 30.32 42.11
Table 6: Comparison of different methods(i.e. smooth v.s. reorder) on LongBench. Group-size(average) is set as 128, key-cache 2bit, value-cache 2bit, window-size 128. We abbreviated PassageRetrieval as PR and MultiFieldQA as MQA.
Appendix E More Results of LongBench Evaluation
We also evaluated LongChat-v1.5-7b-32k and Vicuna-v1.5-7b-16k, which are two famous long-context models fine-tuned based on Llama2-7b. The results in Table7 demonstrate that our SKVQ outperformed previous methods, which highlight the generalizability of our approach.
Model Method LCC RepoBench-P PR-en TREC 2wikimqa GovReport MQA-zh Average
Vicuna-v1.5-7b-16k FP16 51.38 46.18 4.5 69 21.3 27.79 43.74 41.02
RTN 13.22 17.78 1.2 0 0.59 2.39 0.61 8.23
SmoothQuant 40 29.27 1.94 18.25 8.33 14.86 7.19 22.37
RPTQ 40.64 41.4 3.75 59.5 15.92 23.16 19.41 32.68
KIVI 49.32 43.35 5.56 68 23.3 24.47 38.86 39.19
SKVQ 50.98 44.07 6 69 22.04 26.55 40.82 40.20
LongChat-v1.5-7b-32k FP16 54.89 59.05 30.5 66.5 24.58 30.89 35.33 47.27
RTN 5.11 3.73 1.5 0 0.42 0.51 0.08 2.461
SmoothQuant 36.21 31.91 2.45 36.5 13.94 17.21 6.59 24.70
RPTQ 40.4 43.2 8 61 17.31 24.79 20.01 34.01
KIVI 49.86 54.77 20.5 66 23.79 28.75 31.58 43.22
SKVQ 55.01 57.24 22 67 22.4 30.03 31.68 45.37
Table 7: Evaluation resultes of Vicuna and LongChat on LongBench. Group-size(average) 128, key-cache 2bit, value-cache 2bit, window-size 128. We abbreviated PassageRetrieval as PR and MultiFieldQA as MQA. We highlight the result of our method.

### v3

@@ Abstract
Large language models (LLMs) have demonstrated the capability to process extended token sequences, enabling complex tasks such as book comprehension and long-form text generation. However, as context length increases, the key-value (KV) cache required for LLMs consumes substantial memory, becoming a bottleneck for deployment. This paper introduces SKVQ (Sliding-window KV cache Quantization), a strategy designed to address the challenge of extremely low bitwidth KV cache quantization. SKVQ rearranges the channels of the KV cache to enhance channel similarity within quantization groups and applies clipped dynamic quantization at the group level. Furthermore, SKVQ maintains high precision for the most recent window tokens in the KV cache, preserving accuracy for a small yet critical portion of the cache. Our evaluation of LLMs demonstrates that SKVQ achieves high compression ratios while maintaining accuracy, outperforming previous quantization methods. SKVQ enables the quantization of the KV cache to 2-bit keys and 1.5-bit values with minimal accuracy loss. This advancement allows processing context lengths of up to 1M tokens on an 80GB GPU for a 7B parameter model, resulting in up to 7 times faster decoding.
@@ 1 Introduction
Large Language Models (LLMs) have recently demonstrated remarkable success in artificial intelligence. As LLMs advance, the demand for extended context support has increased. For instance, OpenAI GPT-4 Turbo can handle 128k tokens (Achiam et al., 2023), and Google Gemini 1.5 can process up to 1 million tokens (Team et al., 2023). This expanded token capacity enables LLMs to address more complex tasks, including book comprehension, large image analysis, and video processing, enhancing their versatility. LLM inference operates in an auto-regressive manner, generating sentences token by token. To reduce the computation overhead, inference system always store key and value activations in memory known as the key-value (KV) cache, for reuse during subsequent token generation. With the increasing popularity of utilizing LLM for long sequence tasks, the KV cache consumes a significant amount of memory. On the other hand, the large amount of KV cache can also bring a large amount of memory access in the attention mechanism when generating the output tokens. The system will be stuck on the memory access, known as the memory-bound problem in LLM inference (Yuan et al., 2024).
@@
To evaluate the effectiveness of our method, we experiments on models of LLaMA (Touvron et al., 2023) and Mistral (Jiang et al., 2023) family. The experiments show that our methods can quantize the key cache into 2 bits and value cache into 1.5 bits with almost no accuracy drop. Compared with the previous quantization method, our approach can achieve optimal performance under different average bit widths as shown in Figure 1. Our performance analysis shows SKVQ enables 1M context length in a single A100-80GB for a 7b model. As for the inference latency, in the case of batch size 128 and sequence length 200k, the theoretical 7x speedup in decoding phase can be achieved  [^1] The performance analysis can be found in Appendix 9.
@@ 3.1 Clipped Dynamic Quantization with Channel Reorder
To address this problem, some methods have proposed using additional quantization parameters or keeping certain channels in float format to handle outliers (Dettmers et al., 2022). However, we have noticed that the concept of outliers is relative. The channels with the highest values are outliers compared to the medium-sized channels, and the medium-sized channels are outliers compared to the small-sized channels. Other methods propose smoothing the difference between channels by multiplying an extra factor before quantization (Shao et al., 2023; Yue et al., 2024). However, these methods do not take into account the differences in token dimension(sequence length dimension) of KV cache. The magnitude of values can vary between different tokens. We have observed that the variation in magnitude of non-outlier channels is relatively high. Specifically, some channels experience magnitude changes of several times or even dozens of times. Smoothing is not effective in addressing this phenomenon, especially in extremely low bitwidth quantization.
@@
where $P_{k}\in\mathbb{R}^{C_{in}\times C_{in}}$ and $P_{v}\in\mathbb{R}^{C_{in}\times C_{in}}$ are channel reorder matrix of Key and Value respectively  [^2] We fuse the channel reorder index into the the projection weight matrix of attention module. We describe the fusion in the Appendix 6.. In our algorithm, the index is calculated based on the statistical characteristics of each channel. Specifically, we extract the distribution feature of each channel and then use the KMeans algorithm to cluster channels with similar characteristics into the same group. We also compared channel reordering with mathematical equivalent smoothing, and the results in Appendix 10 demonstrated the effectiveness of the former.
@@ 3.2 Sliding Window Quantization Strategy
Locality. Previous research has demonstrated that attention modules exhibit strong locality (Kovaleva et al., 2019; Beltagy et al., 2020; Ge et al., 2023). This locality implies that at each time step, the attention module focuses more on recently generated tokens. We posit that in KV cache quantization, preserving the accuracy of a small but critical portion of the cache is more important than maintaining the larger, less significant content from earlier in the sequence.
@@
Motivated by this observation, we propose a sliding window quantization strategy that maintains high precision for the most recent KV cache of a window of $w$ tokens. The workflow, illustrated in Figure 3, consists of two phases: 1) Prefill phase: For each transformer block, after generating the KV cache, we first compute attention using full-precision KV cache. We then quantize the KV cache, reserving the last $w$ token cache pairs at full precision. 2) Decode phase: We process only the token that slides out of the window at each time step. This approach ensures lossless KV cache generation for each transformer block during the prefill phase. It also enhances the quality of generated content by leveraging the locality of the attention module in the decode phase.
@@
Important KV Cache Filter. Beyond recently generated tokens, certain tokens are particularly sensitive to quantization. We explored additional methods to identify critical tokens whose KV cache should be maintained at high precision. Inspired by (Xiao et al., 2023), we recognized that the initial tokens of a prompt are crucial for the entire generation process. Consequently, we incorporated an attention sink into our filter rules, reserving the first few tokens at high precision. We observed that maintaining a small number of sink tokens at high precision is effective. Given the fixed positions of sink tokens, this approach is easily implementable and was enabled in our experiments. Some cache eviction methods, such as those proposed by (Liu et al., 2023; Zhang et al., 2023), monitor each token’s cumulative attention score, treating these scores as token frequency and retaining only the most frequent tokens (heavy hitters) in the KV cache. While keeping heavy hitters at high precision seems intuitive, we did not implement this approach in our experiments for two reasons: 1) The improvement on prediction accuracy by keeping heavy hitters high precision is not significant. 2) When using FlashAttention (Dao, 2023), directly obtaining attention scores is challenging, making implementation in existing inference frameworks problematic.
We anticipate that superior methods for identifying important KV caches may emerge. Therefore, we have maintained this as an interface (filter rules in Figure 3) in our implementation, allowing for the integration of new filters in future research.
@@
By preserving a small portion of tokens at high precision, we achieve substantial performance gains in long-context tasks while incurring minimal additional overhead. The impact of window size on quantization performance will be discussed in Section 4.3.
@@ 4.2 Main Results and Analysis
[…]
@@
LongBench Results. The performance of SKVQ in the LongBench datasets is summarised in Table 1. We compare our method with SmoothQuant (Xiao et al., 2022), RPTQ (Yuan et al., 2023) KIVI (Liu et al., 2024) and vanilla asymmetric per-token uniform RTN(Round To Nearest) quantization. SmoothQuant and RPTQ are LLM weight-activation quantization schemes. We use them to quantize KV cache without involving model weights and other activation. KIVI (Liu et al., 2024) is a recent 2-bit asymmetric quantization scheme specially designed for KV cache. We set the group size of all the methods to 128. SKVQ utilizes reordering which leads to unequal size of each group. In order to ensure the fairness of the comparison, we control the number of groups in SKVQ to ensure the average group size is 128. The window size in SKVQ is set to 128 and the residual length in KIVI is set to 128. $\alpha$ in SmoothQuant is set to 1.0 to make the smooth transformation completely inclined to KV cache . Table 1 suggests that SKVQ is an effective method for KV cache compression that outperforms previous quantization approaches across various hard long context generation tasks. We also evaluate Vicuna-v1.5-7b-16k and LongChat-v1.5-7b-32k, the results is in Appendix 8.
@@
These results demonstrate that it is practical to quantize the key-value cache into extremely low-bitwidth for these tasks. More result on needle in haystack test can be found in Appendix 7.
@@ 4.3 Ablation Study
Breakdown of different components of SKVQ. We investigate the accuracy impact of various quantization techniques employed in SKVQ. Initially, we utilize RTN and adopt per-token quantization with a group size of 32. Subsequently, we apply other quantization techniques used in SKVQ, including sliding window, clipping, channel reorder, attention sink, and FP8. The LongBench average scores are presented in Table 3. The attention sink size is set to 5, meaning the first 5 token cache pairs are retained at full precision. Our results indicate that sliding-window and channel reorder techniques significantly enhance accuracy. FP8 (E4M3) represents the use of the FP8 datatype to store per-group quantization parameters, such as scale and zero-point. In our study, using FP8 results in a relatively minor accuracy decrease compared to FP16. However, at extremely low-bitwidth and fine-grained group sizes, using FP8 to store quantization parameters significantly reduces the average number of bits. For example, with KV 2-bit quantization and a group size of 32, using FP16 to store quantization parameters results in an average bit count of $2+16*2/32=3$, whereas using FP8 results in an average bit count of $2+82/32=2.5$, which is a 16.7% reduction.
@@ 5 Conclusion
Acknowledgments
We sincerely thank the anonymous COLM reviewers for their valuable comments on this paper. The research is supported under the National Key R&D Program of China (2022ZD0160202).